OpenStudy (agl202):

Quick Question: What is the standard form equation of the line shown below? Graph of a line going through negative 3, negative 1 and 3, 2 y + 1 = one half (x + 3) y = one half x + five halves −x + 2y = 1 x − 2y = −1

OpenStudy (agl202):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (agl202):

I have hard time understanding graphs. :(

jimthompson5910 (jim_thompson5910):

well first off, standard form is Ax+By = C for example, 2x+3y = 6 is in standard form

jimthompson5910 (jim_thompson5910):

so you can cross off 2 choices

OpenStudy (agl202):

The first two?

jimthompson5910 (jim_thompson5910):

yeah you can cross off the first two since they aren't in the form Ax+By = C

OpenStudy (agl202):

So, I think it's D.

OpenStudy (agl202):

right?

jimthompson5910 (jim_thompson5910):

why D ?

OpenStudy (agl202):

Because the x doesn't have negative numbers

jimthompson5910 (jim_thompson5910):

why not plug in each point to see if the equation is satisfied?

jimthompson5910 (jim_thompson5910):

x - 2y = -1 -3 - 2(-1) = -1 ... plug in (x,y) = (-3,-1) simplify both sides. Are the two sides equal?

OpenStudy (agl202):

Oh I see, Makes sense. thx..